6 ( ln l -3) - 4l n l +5 = 11
Answer by Alan3354(69443)

6 ( ln l -3) - 4l n l +5 = 11
---------
If you mean:
6(ln(l) -3) - 4ln(l) +5 = 11
ln(1) = 0
--> 6*(-3) - 4*0 + 5 = 11
-18 + 5 = 11
It doesn't.
I don't know what you have in mind.